If displacement is directly proportional to square of time then prove that the body moves with constant acceleration??

Grade:9

1 Answers

Eshan
5 years ago
Dear student,

We havex\propto t^2\implies x=kt^2

\implies v=\dfrac{dx}{dt}=2kt

\implies a=\dfrac{dv}{dt}=2k=constant

Hope it helps. Regards.